Miller Company’s contribution format income statement for the most recent month is shown below: Total Per Unit Sales (33,000 units) $ 231,000 $ 7.00 Variable expenses 132,000 4.00 Contribution margin 99,000 $ 3.00 Fixed expenses 42,000 Net operating income $ 57,000

Required: (Consider each case independently):

1. What is the revised net operating income if unit sales increase by 12%?

2. What is the revised net operating income if the selling price decreases by $1.10 per unit and the number of units sold increases by 21%?

3. What is the revised net operating income if the selling price increases by $1.10 per unit, fixed expenses increase by $8,000, and the number of units sold decreases by 8%?

4. What is the revised net operating income if the selling price per unit increases by 10%, variable expenses increase by 20 cents per unit, and the number of units sold decreases by 11%?

Can you please provide an explanation of where you get your numbers and what we should multiple or add? Thank you!
Answer #1

All the details regarding computation are provided in the details column which is as below:

1

Total

Per unit

Details

Unit sales (33000 +12% = 36960)

258720

7

36960*7

Variable expenses

147840

4

132000*1.12

Contribution margin

110880

3

Fixed expenses

42000

Net operating income

68880

2

Total

Per unit

Details

Unit sales (33000 +21% = 39930)

235587

7-1.1 = 5.9

39930*5.9

Variable expenses

159720

4

132000*1.21

Contribution margin

75867

1.9

Fixed expenses

42000

Net operating income

33867

3

Total

Per unit

Details

Unit sales (33000 -8% = 30,360 )

245916

7+1.1=8.1

30,360*8.1

Variable expenses

121440

4

132000*0.92

Contribution margin

124476

4.1

Fixed expenses

50,000

42000+8000

Net operating income

74476

4

Total

Per unit

Details

Unit sales (33000 -11% = 29,370 )

226149

7*1.1 = 7.7

29370*7.7

Variable expenses

123354

4.2

29370*4.2

Contribution margin

102795

3.5

Fixed expenses

42,000

42000+8000

Net operating income

60795